# DiCoRe: Enhancing Zero-shot Event Detection via Divergent-Convergent LLM Reasoning

### v1

@@ Abstract
Tanmay Parekh      Kartik Mehta      Ninareh Mehrabi Kai-Wei Chang      Nanyun Peng Computer Science Department, University of California, Los Angeles Amazon AGI Foundations {tparekh, kwchang, violetpeng}@cs.ucla.edu
@@ 1 Introduction
Event Detection (ED) is the task of identifying events by extracting and labeling event triggers Sundheim (1992); Doddington et al. (2004). ED aids in various downstream applications, including news monitoring Tanev et al. (2008), biomedical literature mining Pyysalo et al. (2012), epidemic forecasting Parekh et al. (2024a, b), and legal understanding Francesconi et al. (2010). Training effective ED models requires large amounts of expert-annotated domain-specific data, which is highly costly and labor-intensive. This underlines the need to develop zero-shot systems that can perform ED robustly without using any training data.
@@
We conduct extensive experiments on six datasets from five domains across nine LLMs. Compared with various existing LLM inference works Gao et al. (2023); Wang et al. (2023); Parekh et al. (2025), we show how DiCoRe performs the best with average improvements of 4-5% F1 Trigger Classification and 5.5-6.5% F1 Event Identification over the best baselines. DiCoRe, without any training, also consistently improves over transfer-learning baselines Hsu et al. (2022); Sainz et al. (2023) fine-tuned on 15-30k datapoints by at least 5-12% F1. Furthermore, we demonstrate that DiCoRe provides 1-2% F1 gains while using 15-55x fewer inference tokens relative to strong thinking-based models and chain-of-thought (CoT), highlighting the significance of our proposed divergent-convergent reasoning.
@@ 2 Background and Related Works
Recently, various diverse datasets such as MAVEN Wang et al. (2020), FewEvent Deng et al. (2019), GENEVA Parekh et al. (2023) in general domain, GENIA2011 Kim et al. (2011), GENIA2013 Kim et al. (2013) in biomedical, CASIE Satyapanich et al. (2020) in cybersecurity, PHEE Sun et al. (2022) in pharmacovigilance, SPEED Parekh et al. (2024b), SPEED++ Parekh et al. (2024a) in epidemiology, etc. have been developed. To explore generalizability across these domains/datasets, initial works posed ED as a question-answering Du and Cardie (2020) or machine-reading comprehension problem Liu et al. (2020). Various works explored transfer and joint learning across various IE tasks to build more universal IE models Lu et al. (2022); Fei et al. (2023); Li et al. (2024). Some works have explored posing ED as a generative text-to-text approach with event-based templates Lu et al. (2021); Li et al. (2021); Hsu et al. (2022), even for zero-shot cross-lingual transfer Huang et al. (2022); Parekh et al. (2024b). However, these works require source data training for zero-shot transfer, limiting their utility. Recent works have also explored the utility of zero-shot prompting with LLMs - concluding their sub-par performance Gao et al. (2023); Li et al. (2023a). Other works have explored utilizing LLMs for data generation Ma et al. (2024); Zhang et al. (2024b); Parekh et al. (2025) to aid better generalizability. In our work, we focus on improving LLMs’ zero-shot task generalizability to ED without any model fine-tuning.
@@
LLMs show immense language understanding and generation capabilities, but they need instructions and constraints to aid in meaningful human tasks Ouyang et al. (2022b). However, imposing constraints also reduces LLM reasoning capabilities Tam et al. (2024); Banerjee et al. (2025). To this end, works have explored constrained decoding by altering the output probability distribution Willard and Louf (2023); Netz et al. (2024); Zhang et al. (2024a). Some works explore grammar-aligned decoding strategies Geng et al. (2023); Park et al. (2024). However, such strict enforcement has been shown to hurt LLMs’ generations. Recently, Tam et al. (2024) explored better prompt design on math reasoning to unburden the constraints on the LLM. With similar inspiration, we explore decoupling LLMs from constraints to improve reasoning for IE tasks, specifically Event Detection, in our work.
@@ 3.2 Grounder
Table 1: Main results comparing the zero-shot ED performance of our proposed DiCoRe with all other baselines for the Llama3-8B-Instruct and Llama3-70B-Instruct LLMs. TI: Trigger Identification, TC: Trigger Classification, EI: Event Identification. bold = best performance. (XX) = number of distinct event types.
@@
To reduce the burden of constraint-following on the LLM and ensure strict adherence to the task format, we incorporate a constrained decoding mechanism guided by a finite-state machine (FSM). Inspired by recent work Willard and Louf (2023); Zhang et al. (2024a), the FSM is designed to encode key constraints (Eq. 13) within the decoding process. We construct and demonstrate an FSM to encode constraints for our ED task in Figure 4.
The states of the FSM represent possible decision points (e.g., choice of event type, choice of trigger word, etc.). The state transitions denote possible LLM generations/choices (e.g., list of event types in the ontology $\mathcal{E}$) and are guided by their corresponding LLM generation probabilities. In our FSM in Figure 4, we first decide if there is any event or if it is an event-free sentence (state A). Next, we decide our first event type $e\in\mathcal{E}$ (state B) and then the corresponding trigger $w\in X$ (state C). Finally, we decide if we want to list another event mention or end the predictions (state D). To ensure that the generations are natural, the FSM states are partitioned in alignment with the LLM tokenizer, i.e., the states are chosen such that the sequence of transition tokens is the most probable tokenization of the output text $Y_{g}$.
At each FSM state, we generate using the LLM while constraining its output space by zeroing out the probability mass of all tokens not corresponding to valid state transitions. This enforces that the LLM can only generate tokens permitted by the FSM at each step, effectively guiding the generation process according to the task-specific grammar. As a result, all structural constraints are directly enforced during decoding, ensuring well-formed and ontology-compliant outputs.
@@ 3.3 Judge
4 Experimental Setup
Datasets:
We benchmark our model across six ED datasets spanning five diverse domains, listed as: (1) MAVEN Wang et al. (2020) and (2) FewEvent Deng et al. (2019) from the general domain, (3) ACE Doddington et al. (2004) from the news domain, (4) GENIA Kim et al. (2011), from the biomedical domain, (5) SPEED Parekh et al. (2024b), from the epidemiological/social media domain, (6) CASIE Satyapanich et al. (2020), from the cybersecurity domain. To avoid any distributional biases, following TextEE Huang et al. (2024), we uniformly sample 250 datapoints from the complete dataset for evaluation. [^1] We will release the test splits for reproduction.
@@ 4 Experimental Setup
Table 2: Generalization results for zero-shot ED performance comparing DiCoRe with the best baselines for four other LLMs of Qwen2.5-14B-Instruct, Qwen2.5-72B-Instruct, GPT3.5-turbo, and GPT4o. bold = best performance. (XX) = number of distinct event types.
@@
We consider two major baselines, described below: (1) Multi-event Direct (MD) Gao et al. (2023) directly prompts the LLM to provide the final output in a single pass, and (2) Multi-event Staged (MS) Parekh et al. (2025) decomposes the task into two stages, where the first stage identifies the event and the second stage extracts the corresponding triggers. We also compare with other works like: (3) Binary-event Direct (BD) Lyu et al. (2021); Li et al. (2023c) prompts the LLM to do binary classification for each event, (4) Decompose-Enrich-Extract (DEE) Shiri et al. (2024) utilizes instruction enrichment with schema information for ED, (5) GuidelineEE (GEE) Srivastava et al. (2025), similar to Code4Struct Wang et al. (2023), converts ED into a code-generation problem using Python classes and instantiations, and (6) ChatIE Wei et al. (2023) decomposes ED via multi-turn conversations. We ensure consistent, structured outputs for each baseline to maintain fair comparisons (analysis in § C.1). Furthermore, we add the Judge component to each baseline, if not already present, to ensure robust benchmarking of DiCoRe.
@@
Following Ahn (2006); Parekh et al. (2025) we report the F1 scores for the following three metrics: (1) Trigger Identification (TI) - correct identification of triggers, and (2) Event Identification (EI) - correct classification of event types, and (3) Trigger Classification (TC) - correct identification of the trigger-event pair (event mention). To maintain consistency with traditional span-based evaluations, we used string matching to map the generated outputs to input spans.
@@ 5 Results and Analysis
In this section, we provide our main results and findings, and later provide supporting evidence through our analyses.
@@ 5.1 Main Results
We present the main zero-shot results for all baselines on the six datasets for Llama3 LLMs in Table 1. As seen from the average results (last three columns), DiCoRe performs the best, surpassing the best baseline of multi-event staged (MS) by a significant margin of 5.5-8% TI, 4-8.5% EI, and 4-5% TC. The performance disparity across different task decomposition methods of ChatIE, MS, and DiCoRe highlights how our divergent-convergent decomposition of Dreamer-Grounder provides a stronger inductive bias. Other baselines perform relatively better on the fewer-event and simpler datasets like GENIA/SPEED, but DiCoRe shows strong dominance on the larger event datasets like MAVEN/FewEvent/ACE.
Generalization across LLMs:
@@
5.2 Comparison with Fine-tuned Transfer-learning Methods
@@
Sentence Best Baseline Dreamer Grounder Judge
Prediction Prediction Prediction Prediction
@@ Appendix B Additional Experimental Details
In § 4, we provided brief details about our experimental and implementation details. Here, we provide more intricate details about our different experimental setups and baseline models.
Dataset Domain # Doc # Event Avg. Doc
Mentions Length
MAVEN General 250 623 24.5
FewEvent General 250 250 30.5
ACE News 250 71 13.2
GENIA Biomedical 250 2472 251.3
SPEED Epidemiology 250 258 32.4
CASIE Cybersecurity 50 291 283.1
B.1 Dataset Statistics
Our experimental setup is a pure zero-shot setup where we do not use any training data. We provide statistics about the evaluation splits of the different datasets in Table 7. We follow TextEE Huang et al. (2024) for the evaluation setup and consider a uniform random split of 250 test samples from each dataset to avoid any train-test split bias. Since CASIE is a smaller dataset, we only use 50 test samples for this dataset. The table highlights the domain diversity of the datasets covering common domains like news and general, while also focusing on technical domains like biomedical and epidemiology. The datasets also show variation in the density, with ACE, FewEvent, and SPEED being sparse with  1 event mention/sentence. On the other hand, MAVEN, CASIE, and GENIA are denser with 2.5-10 event mentions/passage. Finally, we also show the variation in token length, with ACE being the lowest with 13 average tokens, while GENIA and CASIE are longer with 250-280 average tokens per document.
B.2 Additional Implementation Details
In this section, we provide additional implementation details for DiCoRe and the various baselines. For open-source models, we ran them locally on NVIDIA RTX A6000/A100 machines with support for 8 GPUs.
B.2.1 DiCoRe
@@
B.2.2 Multi-event Direct (MD)
@@
B.2.3 Multi-event Staged (MS)
Multi-event staged (MS) Parekh et al. (2025) was introduced as a way of forward generation to ensure higher trigger quality. We extend that in our work to build a strong task decomposition baseline. Simply, this model first extracts the event types from the texts in Stage 1 and then extracts triggers specific to these event types in Stage 2. We provide an illustration of the two stages of MS in Figures 9 and  10. In this case, the first stage majorly only focuses on the event-specific constraints, while the second stage is focused on the trigger-specific ones.
@@ B.2.3 Multi-event Staged (MS)
B.2.4 Binary-event Direct (BD)
Binary-event direct (BD) Lyu et al. (2021); Li et al. (2023c) has been a popular paradigm pre-dating LLMs when smaller generative text-to-text models were used. It drastically reduces the LLM’s constraints by making the LLM focus on a single event type at a time, i.e., it prompts the LLM in a multi-event direct manner, but for each event type separately. Finally, the predictions are aggregated and output as the final prediction. We provide an illustration of the prompt in Figure 11. Overall, this is a highly expensive method, especially for larger event datasets.
@@ B.2.4 Binary-event Direct (BD)
B.2.5 Decompose-Enrich-Extract (DEE)
@@ C.1 Structured v/s Unstructured Output
C.2 Complete Results for Transfer Learning Baselines

### v2

@@ Abstract
Tanmay Parekh Kartik Mehta Ninareh Mehrabi§thanks: Work done while at Amazon. Kai-Wei Chang Nanyun Peng Computer Science Department, University of California, Los Angeles Amazon AGI Foundations     §Meta {tparekh, kwchang, violetpeng}@cs.ucla.edu
@@ 1 Introduction
Event Detection (ED) is the task of identifying events by extracting and labeling event triggers Sundheim (1992); Doddington et al. (2004). ED aids in various downstream applications, including news monitoring Tanev et al. (2008), biomedical literature mining Pyysalo et al. (2012), epidemic forecasting Parekh et al. (2024b, c), and legal understanding Francesconi et al. (2010). Training effective ED models requires large amounts of expert-annotated domain-specific data, which is highly costly and labor-intensive. This underlines the need to develop zero-shot systems that can perform ED robustly without using any training data.
@@
We conduct extensive experiments on six datasets from five domains across nine LLMs. Compared with various existing LLM inference works Gao et al. (2023); Wang et al. (2023); Parekh et al. (2025a), we show how DiCoRe performs the best with average improvements of 4-5% F1 Trigger Classification and 5.5-6.5% F1 Event Identification over the best baselines. DiCoRe, without any training, also consistently improves over transfer-learning baselines Hsu et al. (2022); Sainz et al. (2023) fine-tuned on 15-30k datapoints by at least 5-12% F1. Furthermore, we demonstrate that DiCoRe provides 1-2% F1 gains while using 15-55x fewer inference tokens relative to strong thinking-based models and chain-of-thought (CoT), highlighting the significance of our proposed divergent-convergent reasoning.
@@ 2 Background and Related Works
Recently, various diverse datasets such as MAVEN Wang et al. (2020), FewEvent Deng et al. (2019), GENEVA Parekh et al. (2023) in general domain, GENIA2011 Kim et al. (2011), GENIA2013 Kim et al. (2013) in biomedical, CASIE Satyapanich et al. (2020) in cybersecurity, PHEE Sun et al. (2022) in pharmacovigilance, SPEED Parekh et al. (2024c), SPEED++ Parekh et al. (2024b) in epidemiology, etc. have been developed. To explore generalizability across these domains/datasets, initial works posed ED as a question-answering Du and Cardie (2020) or machine-reading comprehension problem Liu et al. (2020). Various works explored transfer and joint learning across various IE tasks to build more universal IE models Lu et al. (2022); Fei et al. (2023); Li et al. (2024). Some works have explored posing ED as a generative text-to-text approach with event-based templates Lu et al. (2021); Li et al. (2021); Hsu et al. (2022), even for zero-shot cross-lingual transfer Huang et al. (2022); Parekh et al. (2024a). However, these works require source data training for zero-shot transfer, limiting their utility. Recent works have also explored the utility of zero-shot prompting with LLMs - concluding their sub-par performance Gao et al. (2023); Li et al. (2023a). Other works have explored utilizing LLMs for data generation Ma et al. (2024); Zhang et al. (2024b); Parekh et al. (2025a) to aid better generalizability. In our work, we focus on improving LLMs’ zero-shot task generalizability to ED without any model fine-tuning.
@@
LLMs show immense language understanding and generation capabilities, but they need instructions and constraints to aid in meaningful human tasks Ouyang et al. (2022b). However, imposing constraints also reduces LLM reasoning capabilities Tam et al. (2024); Tian et al. (2024); Banerjee et al. (2025). To this end, works have explored constrained decoding by altering the output probability distribution Willard and Louf (2023); Netz et al. (2024); Zhang et al. (2024a). Some works explore grammar-aligned decoding strategies Geng et al. (2023); Park et al. (2024). However, such strict enforcement has been shown to hurt LLMs’ generations. Recently, Tam et al. (2024) explored better prompt design on math reasoning to unburden the constraints on the LLM. With similar inspiration, we explore decoupling LLMs from constraints to improve reasoning in our work. Although, we only explore the task of Event Detection, we believe our work could benefit other structured tasks in Information Extraction Li et al. (2023c); Wang et al. (2025), Document Understanding Suvarna et al. (2024), Question Answering Rajpurkar et al. (2016); Parekh et al. (2025b), and Dialogue Generation Parekh et al. (2020); Chen et al. (2020).
@@ 3.2 Grounder
Table 1: Main results comparing the zero-shot ED performance of our proposed DiCoRe with all other baselines for the Llama3-8B-Instruct and Llama3-70B-Instruct LLMs. TI: Trigger Identification, TC: Trigger Classification, EI: Event Identification. bold = best performance. ($X$) = number of distinct event types.
@@
To reduce the burden of constraint-following on the LLM and ensure strict adherence to the task format, we incorporate a constrained decoding mechanism guided by a finite-state machine (FSM). Inspired by recent work Willard and Louf (2023); Zhang et al. (2024a), the FSM explicitly encodes structural and task-specific constraints (Eq. 13) within the decoding process. We construct and demonstrate an FSM to encode constraints for our ED task in Figure 4.
The FSM states represent decision points (e.g., whether the sentence contains an event, which event type $e\in\mathcal{E}$ to choose, which trigger $w\in X$ to assign, etc.), and the transitions denote valid LLM generations at each point (e.g., list of event types in $\mathcal{E}$, trigger words in the sentence). As shown in Figure 4, generation proceeds step by step: starting with the event/no-event decision (state A), followed by selecting an event type (state B), then its trigger (state C), and finally deciding whether to generate another event mention or terminate (state D). To ensure that the generations are natural, the FSM states are partitioned in alignment with the LLM tokenizer, i.e., the states are chosen such that the sequence of transition tokens is the most probable tokenization of the output text $Y_{g}$.
We implement this FSM using the Outlines library Willard and Louf (2023) integrated into a vLLM inference framework Kwon et al. (2023). The module takes as input the ontology, input sentence, LLM, and output JSON schema (potentially expressed as a grammar). Each FSM state transition is encoded as an Outlines choices list, thereby restricting the LLM’s output vocabulary to only valid strings for that transition. For example, the set of possible event types or candidate trigger words is directly provided as the restricted vocabulary, and transitions with a single option are handled deterministically. The selected string then determines the next FSM state.
This design enforces structural validity during decoding: at each step, tokens not corresponding to valid FSM transitions are zeroed out, ensuring the LLM can only generate ontology-compliant outputs. Our implementation currently supports generation of JSON tuples of the form (event type, trigger), making it directly applicable to any ED dataset. More generally, because the transition and state mappings can be automatically constructed from the grammar of task constraints, the approach is customizable to other output formats and structured prediction tasks.
@@ 3.3 Judge
Table 2: Generalization results for zero-shot ED performance comparing DiCoRe with the best baselines for four other LLMs of Qwen2.5-14B-Instruct, Qwen2.5-72B-Instruct, GPT3.5-turbo, and GPT4o. bold = best performance. ($X$) = number of distinct event types.
4 Experimental Setup
Datasets:
We benchmark our model across six ED datasets spanning five diverse domains, listed as: (1) MAVEN Wang et al. (2020) and (2) FewEvent Deng et al. (2019) from the general domain, (3) ACE Doddington et al. (2004) from the news domain, (4) GENIA Kim et al. (2011), from the biomedical domain, (5) SPEED Parekh et al. (2024c), from the epidemiological/social media domain, (6) CASIE Satyapanich et al. (2020), from the cybersecurity domain.
Dataset Domain # Doc # Event Avg. Doc
Mentions Length
MAVEN General 250 623 24.5
FewEvent General 250 250 30.5
ACE News 250 71 13.2
GENIA Biomedical 250 2472 251.3
SPEED Epidemiology 250 258 32.4
CASIE Cybersecurity 50 291 283.1
We provide statistics about the test splits of the different datasets in Table 3. To avoid any distributional biases, following TextEE Huang et al. (2024), we uniformly sample 250 datapoints from the combined train-dev-test splits of each dataset for evaluation. Since CASIE is a smaller dataset, we only use 50 test samples for this dataset. The table highlights the domain diversity of the datasets covering common domains like news and general, while also focusing on technical domains like biomedical and epidemiology. The datasets also show variation in the density, with ACE, FewEvent, and SPEED being sparse with upto 1 event mention/sentence. On the other hand, MAVEN, CASIE, and GENIA are denser with 2.5-10 event mentions/passage. Finally, we also show the variation in token length, with ACE being the lowest with an average of 13 tokens, while GENIA and CASIE are longer with 250-280 average tokens per input document.
@@ 4 Experimental Setup
We consider two major baselines, described below: (1) Multi-event Direct (MD) Gao et al. (2023) directly prompts the LLM to provide the final output in a single pass, and (2) Multi-event Staged (MS) Parekh et al. (2025a) decomposes the task into two stages, where the first stage identifies the event and the second stage extracts the corresponding triggers. We also compare with other works like: (3) Binary-event Direct (BD) Lyu et al. (2021); Li et al. (2023d) prompts the LLM to do binary classification for each event, (4) Decompose-Enrich-Extract (DEE) Shiri et al. (2024) utilizes instruction enrichment with schema information for ED, (5) GuidelineEE (GEE) Srivastava et al. (2025), similar to Code4Struct Wang et al. (2023), converts ED into a code-generation problem using Python classes and instantiations, and (6) ChatIE Wei et al. (2023) decomposes ED via multi-turn conversations. We ensure consistent, structured outputs for each baseline to maintain fair comparisons (analysis in § C.1). Furthermore, we add the Judge component to each baseline, if not already present, to ensure robust benchmarking of DiCoRe.
@@
Following Ahn (2006); Parekh et al. (2025a) we report the F1 scores for the following three metrics: (1) Trigger Identification (TI) - correct identification of triggers, and (2) Event Identification (EI) - correct classification of event types, and (3) Trigger Classification (TC) - correct identification of the trigger-event pair (event mention). To maintain consistency with traditional span-based evaluations, we used string matching to map the generated outputs to input spans.
@@ 5 Results and Analysis
In this section, we provide our main results and findings, and later provide supporting evidence through our analyses. We also provide additional experimental results and error anlaysis in the Appendix (§ C).
@@ 5.1 Main Results
We present the main zero-shot results for all baselines on the six datasets for Llama3 LLMs in Table 1. As seen from the average results (last three columns), DiCoRe performs the best, surpassing the best baseline of multi-event staged (MS) by a significant margin of 5.5-8% TI, 4-8.5% EI, and 4-5% TC. The performance disparity across different task decomposition methods of ChatIE, MS, and DiCoRe highlights how our divergent-convergent decomposition of Dreamer-Grounder provides a stronger inductive bias. Other baselines perform relatively better on datasets like GENIA/SPEED, as these are simpler datasets with fewer event types; thus, requiring lesser cognitive reasoning. However, on the high-event datasets like MAVEN/FewEvent/ACE which require more complex reasoning, DiCoRe with its divergent-convergent reasoning shows more significant improvement over the baselines.
@@
Generalization across LLMs:
@@
Sentence Best Baseline Dreamer Grounder Judge
Prediction Prediction Prediction Prediction
5.2 Comparison with Fine-tuned Transfer-learning Methods
@@ 6 Conclusion and Future Work
Acknowledgments
We express our gratitude to Po-Nien Kung and Haoyi Qiu for their valuable time, reviews of our work, and constructive feedback. We thank the anonymous reviewers, area chairs, and the program committee for their reviews and feedback. We are also grateful to Amazon for supporting this work through the Amazon Science Ph.D. Fellowship awarded to Tanmay Parekh. This work was also partially supported by the National Science Foundation CAREER award #2339766 and the Amazon AGI Research Award awarded to Nanyun Peng.
@@ Appendix B Additional Experimental Details
In § 4, we provided brief details about our experimental and implementation details. Here, we provide additional implementation details for DiCoRe and the various baselines. For open-source models, we ran them locally on NVIDIA RTX A6000/A100 machines with support for 8 GPUs.
B.1 DiCoRe
@@
B.2 Multi-event Direct (MD)
@@
B.3 Multi-event Staged (MS)
Multi-event staged (MS) Parekh et al. (2025a) was introduced as a way of forward generation to ensure higher trigger quality. We extend that in our work to build a strong task decomposition baseline. Simply, this model first extracts the event types from the texts in Stage 1 and then extracts triggers specific to these event types in Stage 2. We provide an illustration of the two stages of MS in Figures 9 and  10. In this case, the first stage majorly only focuses on the event-specific constraints, while the second stage is focused on the trigger-specific ones.
@@ B.3 Multi-event Staged (MS)
B.4 Binary-event Direct (BD)
Binary-event direct (BD) Lyu et al. (2021); Li et al. (2023d) has been a popular paradigm pre-dating LLMs when smaller generative text-to-text models were used. It drastically reduces the LLM’s constraints by making the LLM focus on a single event type at a time, i.e., it prompts the LLM in a multi-event direct manner, but for each event type separately. Finally, the predictions are aggregated and output as the final prediction. We provide an illustration of the prompt in Figure 11. Overall, this is a highly expensive method, especially for larger event datasets.
@@ B.4 Binary-event Direct (BD)
B.5 Decompose-Enrich-Extract (DEE)
@@ C.1 Structured v/s Unstructured Output
Prompt MAVEN (168) FewEvent (100) ACE (33)
Style TI TC EI TI TC EI TI TC EI
ChatIE 33.7 ($\pm$ 0.9) 7.3 ($\pm$ 0.6) 13.8 ($\pm$ 0.6) 20.8 ($\pm$ 0.8) 10.2 ($\pm$ 0.6) 27.6 ($\pm$ 0.4) 30.6 ($\pm$ 1.4) 24.9 ($\pm$ 0.9) 46.8 ($\pm$ 0.9)
GEE 19.1 ($\pm$ 1.7) 1.9 ($\pm$ 0.7) 6.8 ($\pm$ 0.6) 11.7 ($\pm$ 1.5) 5.9 ($\pm$ 1.7) 14.0 ($\pm$ 1.7) 30.0 ($\pm$ 1.7) 21.3 ($\pm$ 0.7) 27.4 ($\pm$ 1.3)
DEE 33.7 ($\pm$ 1.4) 6.0 ($\pm$ 0.7) 9.2 ($\pm$ 0.4) 21.1 ($\pm$ 0.5) 10.6 ($\pm$ 0.4) 17.8 ($\pm$ 0.2) 26.9 ($\pm$ 0.8) 19.8 ($\pm$ 0.7) 36.1 ($\pm$ 0.8)
BD 54.5 ($\pm$ 0.6) 10.7 ($\pm$ 0.7) 12.3 ($\pm$ 0.5) 22.3 ($\pm$ 1.7) 9.9 ($\pm$ 0.9) 15.0 ($\pm$ 0.8) 34.2 ($\pm$ 1.9) 19.5 ($\pm$ 1.5) 31.4 ($\pm$ 1.1)
MD 45.9 ($\pm$ 1.2) 2.8 ($\pm$ 0.2) 4.0 ($\pm$ 0.3) 25.2 ($\pm$ 0.7) 9.5 ($\pm$ 0.2) 15.2 ($\pm$ 0.6) 35.6 ($\pm$ 1.2) 22.4 ($\pm$ 0.8) 30.1 ($\pm$ 0.5)
MS 46.2 ($\pm$ 1.3) 10.3 ($\pm$ 0.7) 11.2 ($\pm$ 0.8) 20.2 ($\pm$ 1.1) 10.2($\pm$ 0.7) 17.0 ($\pm$ 1.1) 26.7 ($\pm$ 1.4) 17.6 ($\pm$ 0.6) 23.1 ($\pm$ 0.9)
DiCoRe 53.5 ($\pm$ 1.1) 14.4 ($\pm$ 0.7) 17.4 ($\pm$ 0.6) 26.1 ($\pm$ 0.4) 15.7 ($\pm$ 0.7) 25.0 ($\pm$ 0.6) 40.3 ($\pm$ 1.9) 36.3 ($\pm$ 1.2) 47.9 ($\pm$ 0.8)
Table 10: Main results along with error bars indicating confidence intervals for the zero-shot ED performance of our proposed DiCoRe with all other baselines for the Llama3-8B-Instruct. TI: Trigger Identification, TC: Trigger Classification, EI: Event Identification. bold = best performance. ($X$) = number of distinct event types.
C.2 Statistical Significant Testing
To verify that our results are statistically significant, we provide error bars indicating confidence intervals in Table 10 for MAVEN, FewEvent, and ACE datasets using the Llama3-8B-Instruct. These results demonstrate how our experimental improvements are statistically sound. We also test and demonstrate that the improvements by DiCoRe are statistically significant (t-test using $p<0.01$).
Prompt Style ACE MAVEN
TI TC EI TI TC EI
MD 36.4 28.6 34.7 45.9 3.1 4.4
MS 28.2 20.7 25.2 45.6 11.1 12.7
DiCoRe 47.2 38.3 48.3 53.2 15.3 18.7
C.3 Results on larger test data size
Our experimental data comprised 250 samples for evaluation to keep our findings/results consistent with the the previous work of TextEE Huang et al. (2024). Here, we provide additional experiments on larger test data size of 1000 samples for ACE and MAVEN datasets in Table 11. Similar to patterns in the main results, DiCoRe outperforms the baselines with gains upto 14% F1 for ACE and 8% F1 for MAVEN.
C.4 Complete Results for Transfer Learning Baselines